OpenStudy (anonymous):

please check alg 2: What are the discontinuities of the function f(x) = x^2-36/4x+24? --> x can not = -6 correct? What is the oblique asymptote of the function f(x) = x^2+5x+6/x-4? --> y =x+9 correct? What are the vertical asymptotes of the function f(x) = 5x+1/x^2+x-6? --> x can not = -2, x = 3 correct? What are the zeros of the function f(x) = 3x^2-3x-6/3x-6? --> 2 correct?

OpenStudy (anonymous):

need help: Consider the function f(x) = x^2-16/x+3. Determine all asymptotes of this function including horizontal, vertical, and oblique (slant).

OpenStudy (anonymous):

please check alg 2: Consider the function f(x) = x^2+x-12/x-3. Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph. disc = -4? x int = (3,-4)? y int = 3? basic shape = ellipse?